Tim and Lynne are jogging in opposite directions. They start at the same place and at the same time. Time jogs at a rate of 5 1/2 miles per hour, and Lynne jogs at a rate of 4 3/4 miles per hour. How far apart will they be in 1 1/4 hours?
* I know how to do the problem but don't know how to set it up*

Answer by alicealc(293)

Tim's distance will be:
5 1/2 * 1 1/4 = 11/2 * 5/4 = 55/8 miles

Lynne's distance will be:
4 3/4 * 1 1/4 = 19/4 * 5/4 = 95/16 miles

so, they will be:
55/8 + 95/16 = 110/16 + 95/16 = 205/16 = 12 13/16 miles apart to each other